OpenStudy (18jonea):

A quadratic equation factored to 0 = (x-4)(x+3) would have solutions of x = -4 and x = 3. True False is this false @Directrix

Directrix (directrix):

0 = (x-4)(x+3) Use the Zero Product Property x - 4 = 0 or x + 3 = 0 x = ? or x = ?

Directrix (directrix):

>>is this false Correct. It is false.

OpenStudy (18jonea):

olve 6x2 = 48. If necessary, round your answer to the nearest hundredth. How would I do this

OpenStudy (18jonea):

@Directrix

Directrix (directrix):

6x2 = 48 Divide both sides by 6. x^2 = 8 Find the square root of 8. x = + ? or - ?

OpenStudy (18jonea):

2.83

OpenStudy (18jonea):

@Directrix